# Explaining latent representations of generative models with large multimodal models

### v1

@@ Abstract
Learning interpretable representations of data generative latent factors is an important topic for the development of artificial intelligence. With the rise of the large multimodal model, it can align images with text to generate answers. In this work, we propose a framework to comprehensively explain each latent factor in the generative models using a large multimodal model. We further measure the uncertainty of our generated explanations, quantitatively evaluate the performance of explanation generation among multiple large multimodal models, and qualitatively visualize the variations of each latent factor to learn the disentanglement effects of different generative models on explanations. Finally, we discuss the explanatory capabilities and limitations of state-of-the-art large multimodal models.
@@ 1 Introduction
Recently, one of the most powerful LMMs is the instruction-following LMM (Li, 2023). LLaVA and InstructBLIP are two instruction-following LMMs that achieve SOTA performance on many datasets. Instruction-following models use instruction tuning to enhance their abilities to understand and follow human-given instructions. Instruction tuning involves further fine tuning LLMs using $\langle instruction,response\rangle$ pairs to better align human intent with model behavior (Wang et al., 2024). LLaVA was introduced in the paper Visual Instruction Tuning (Liu et al., 2023b), and then further improved in Improved Baselines with Visual Instruction Tuning (referred to as LLaVA-1.5) (Liu et al., 2023a). Likewise, InstructBLIP (Dai et al., 2023) is a large multimodal model that adds instruction tuning on the basis of its previous version BLIP-2 (Li et al., 2023). Google Bard is a conversational AI service developed by Google, initially powered by LaMDA with a range of models to follow. In this work, we propose a framework to comprehensively explain each latent factor in generative models and evaluate the performance of explanation generation of GPT-4-vision with several popular LMMs: Google Bard, LLaVA-1.5, and InstructBLIP. To the best of our knowledge, we are the first to use LMM to explain the latent representations of the generative models.
@@ 2 Methods
Problem Formulation We start from a dataset $\mathcal{D}=\{X,Z\}$ where the images $\mathbf{x}\in\mathbb{R}^{N}$ and the data generative latent factors $\mathbf{z}\in\mathbb{R}^{M}$. We then can train a generative model that learns the joint distribution of the data $\mathbf{x}$ and latent factors $\mathbf{z}$. Here, our goal is to explain each latent factor ${z}_{i}$ individually, where i ranges from 1 to the number of latent dimensions, and quantify the uncertainty of the explanation to make sure the explanations presented are reliable and responsible.
@@
In this work, we confront two primary challenges associated with the representation and interpretation of the latent space. Firstly, the latent space is difficult to explicitly represent. To address this, we interpolate a specific latent dimension ${z}_{i}$ at a time and subsequently decode these into an image sequence. This method allows us to visualize the latent representations. Secondly, not all latent factors have semantic meaning. To tackle this issue, we introduce an uncertainty approach to distinguish between interpretable and uninterpretable latent factors.
@@
Figure 1: The model framework consists of generating an image sequence with a progressively varying latent factor, combining it with a prompt to pass to a large multimodal model to provide some response samples and finally utilizing an uncertainty measure to select an explanation for that specific latent factor and decide whether there is a clear explanation to display.
@@
Framework To start with, we trained three generative models for each dataset. After the generative models are trained, a latent representation ${\mathbf{z}}$ is first sampled from an isotropic unit Gaussian distribution $\mathcal{N}(\mathbf{0},{\bm{I}})$. Given a certain latent factor ${z}_{i}$, we perturb ${z}_{i}$ to observe the possible value range of ${z}_{i}$ and keep the other latent factors constant. A sequence of images can then be generated by decoding a series of manipulated latent vectors. As a result, this generated sequence of images shows how a latent factor ${z}_{i}$ changes gradually. Further, we pass this image sequence with a prompt to a large multimodal model to explain the latent factor, as well as its changing pattern. Finally, we quantify the uncertainty to decide which response to select as an explanation and whether to display the selected explanation as shown in Figure 1.
@@
Not every latent factor is interpretable, so we need to find an approach to determine whether a latent factor is interpretable. Similarity and entropy are two major ways to measure uncertainty in natural language generation(NLG). Since the likelihood of tokens is not available in GPT-4-vision, we used the measure of similarity in this paper. However, the use of similarity is not a limitation of the uncertainty measure. Other methods, such as predictive entropy, semantic entropy (Kuhn et al., 2023), and P(true) (Kadavath et al., 2022), could also be used to quantify uncertainty if applicable. To measure the uncertainty of the responses of our large multimodal model, GPT-4-vision, we sampled $n$ times from the GPT-4-vision to generate the responses $R=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average similarity of the responses $R:\frac{1}{C}\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$, where $C=n*(n-1)/2$. To find the threshold of the interpretability of the latent factors, we denote the true label of the interpretability of each latent factor as 1 if a human can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for each row of generated images is its predicted probability. We can then compute their AUC (area under the ROC curve) based on different thresholds and choose the one with the highest AUC as our threshold $\varepsilon$. Our final output explanation is the explanation that has the highest mean pairwise similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, we will output there is no clear explanation.
@@ 3 Experiments
Visual pattern generation We train three representative variational autoencoder(VAE) models, the standard VAE (Kingma & Welling, 2014; Rezende et al., 2014), $\beta$-VAE (Higgins et al., 2016), and $\beta$-TCVAE (Chen et al., 2018), with three aforementioned datasets. For each trained model, we manipulate one latent factor between [-3, 3] at a time while keeping others constant, and the trained decoder of the model can produce a series of images that reflect variations along that specific latent dimension. This process is repeated for each latent factor, resulting in $m\times k$ images, where $m$ is the number of latent factors and $k$ is the number of assigned values for each latent factor. We set 6 latent factors and assign the values as torch.range(-3, 3, 0.6), so $m=6$ and $k=11$ here.
Explanation generation We use GPT-4-vision as our explanation generator and compare it with a wide range of other large multimodal models: Google Bard, LLaVA-v1.5-13b (Liu et al., 2023a), and InstructBLIP (Dai et al., 2023). For GPT-4-vision, LLaVA-v1.5-13b, and InstructBLIP, we all set $temperature=1,top\_p=1$. We then pass the generated images of each latent factor along with a prompt to the LLMs to produce an explanation for each latent factor. For human annotations, two annotators provide two explanations with various sentence expressions for each image sequence, so we have four human explanations for each image sequence as references. Furthermore, We evaluate the explanations across these LMMs with human annotations using BLEU(Papineni et al., 2002), ROUGE-L(Lin, 2004), METEOR(Banerjee & Lavie, 2005), and BERTScore F1(Zhang et al., 2019).
@@
lexical similarity
0.6898
0.9600
0.9412
0.9796
cosine similarity
0.9694 0.9684 1.0
0.9388
@@ 5 Case study
Figure 2 shows the explanations generated by our framework for the 3dshapes dataset. If the certainty score of the generated explanation is equal or greater than the certainty threshold ($\varepsilon=0.7434$), it will display the explanation selected. Otherwise, it will return “No clear explanation” like ${z}_{1}$ and ${z}_{5}$. Our framework can generate an explanation for each latent factor if the latent factor has an evident pattern. The explanation not only can tell what the latent factor is, but also can illustrate how the latent factor changes. Sometimes, the identified latent factor can be wrong, such as it misclassifies the wall orientation as background color in ${z}_{3}$. This is due to the deficiency in the LMM’s visual capabilities and will be elaborated in the later limitation section.
@@ 7 Conclusion
In this work, we propose a framework to comprehensively explain each latent factor in the generative models and visualize the variations of each latent factor. We first analyze the certainty scores of the explanations. The certainty score can tell us if there is a clear trend in the latent factor. We also evaluate and compare the generated explanations with various LMMs. The result shows that GPT-4-vision outperforms other large multimodal models, and we further discuss its limitations and the explanatory capabilities of various LMMs. At last, we evaluate the effects of latent factor disentanglement on the generated explanations. We believe our approach provides an efficient, explainable, and reliable way to learn the latent representations of generative models.

### v2

@@ Abstract
Learning interpretable representations of data generative latent factors is an important topic for the development of artificial intelligence. With the rise of the large multimodal model, it can align images with text to generate answers. In this work, we propose a framework to comprehensively explain each latent variable in the generative models using a large multimodal model. We further measure the uncertainty of our generated explanations, quantitatively evaluate the performance of explanation generation among multiple large multimodal models, and qualitatively visualize the variations of each latent variable to learn the disentanglement effects of different generative models on explanations. Finally, we discuss the explanatory capabilities and limitations of state-of-the-art large multimodal models.
@@ 1 Introduction
Recently, one of the most powerful LMMs is the instruction-following LMM (Li, 2023). LLaVA and InstructBLIP are two instruction-following LMMs that achieve SOTA performance on many datasets. Instruction-following models use instruction tuning to enhance their abilities to understand and follow human-given instructions. Instruction tuning involves further fine tuning LLMs using $\langle instruction,response\rangle$ pairs to better align human intent with model behavior (Wang et al., 2024). LLaVA was introduced in the paper Visual Instruction Tuning (Liu et al., 2024), and then further improved in Improved Baselines with Visual Instruction Tuning (referred to as LLaVA-1.5) (Liu et al., 2023). Likewise, InstructBLIP (Dai et al., 2024) is a large multimodal model that adds instruction tuning on the basis of its previous version BLIP-2 (Li et al., 2023). Google Bard is a conversational AI service developed by Google, initially powered by LaMDA with a range of models to follow. In this work, we propose a framework to comprehensively explain each latent variable in generative models and evaluate the performance of explanation generation of GPT-4-vision with several popular LMMs: Google Bard, LLaVA-1.5, and InstructBLIP. To the best of our knowledge, we are the first to use LMM to explain the latent representations of the generative models.
@@ 2 Methods
Problem Formulation We start from a dataset $\mathcal{D}=\{X,Z\}$ where the images $\mathbf{x}\in\mathbb{R}^{N}$ and the data generative latent variables $\mathbf{z}\in\mathbb{R}^{M}$. We then can train a generative model that learns the joint distribution of the data $\mathbf{x}$ and latent variables $\mathbf{z}$. Here, our goal is to explain each latent variable ${z}_{i}$ individually, where i ranges from 1 to the number of latent dimensions, and quantify the uncertainty of the explanation to make sure the explanations presented are reliable and responsible.
@@
In this work, we confront two primary challenges associated with the representation and interpretation of the latent space. Firstly, the latent space is difficult to explicitly represent. To address this, we interpolate a specific latent dimension ${z}_{i}$ at a time and subsequently decode these into an image sequence. This method allows us to visualize the latent representations. Secondly, not all latent variables have semantic meaning. To tackle this issue, we introduce an uncertainty approach to distinguish between interpretable and uninterpretable latent variables.
@@
Figure 1: The model framework consists of generating an image sequence with a progressively varying latent variable, combining it with a prompt to pass to a large multimodal model to provide some response samples and finally utilizing an uncertainty measure to select an explanation for that specific latent variable and decide whether there is a clear explanation to display.
@@
Framework To start with, we trained three generative models for each dataset. After the generative models are trained, a latent representation ${\mathbf{z}}$ is first sampled from an isotropic unit Gaussian distribution $\mathcal{N}(\mathbf{0},{\bm{I}})$. Given a certain latent variable ${z}_{i}$, we perturb ${z}_{i}$ to observe the possible value range of ${z}_{i}$ and keep the other latent variables constant. A sequence of images can then be generated by decoding a series of manipulated latent vectors. As a result, this generated sequence of images shows how a latent variable ${z}_{i}$ changes gradually. Further, we pass this image sequence with a prompt to a large multimodal model to explain the latent variable, as well as its changing pattern. Finally, we quantify the uncertainty to decide which response to select as an explanation and whether to display the selected explanation as shown in Figure 1.
@@
Not every latent variable is interpretable, so we need to find an approach to determine whether a latent variable is interpretable. Similarity and entropy are two major ways to measure uncertainty in natural language generation(NLG). Since the likelihood of tokens is not available in GPT-4-vision, we used the measure of similarity in this paper. However, the use of similarity is not a limitation of the uncertainty measure. Other methods, such as predictive entropy, semantic entropy (Kuhn et al., 2023), and P(true) (Kadavath et al., 2022), could also be used to quantify uncertainty if applicable. To measure the uncertainty of the responses of our large multimodal model, GPT-4-vision, we sampled $n$ times from the GPT-4-vision to generate the responses $R=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average similarity of the responses $R:\frac{1}{C}\sum_{i=1}^{n}\sum_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$, where $C=n*(n-1)/2$. To find the threshold of the interpretability of the latent variables, we denote the true label of the interpretability of each latent variable as 1 if a human can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for each row of generated images is its predicted probability. We can then compute their AUC (area under the ROC curve) based on different thresholds and choose the one with the highest AUC as our threshold $\varepsilon$. Our final output explanation is the explanation that has the highest mean pairwise similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, we will output there is no clear explanation.
@@ 3 Experiments
Visual pattern generation We train three representative variational autoencoder(VAE) models, the standard VAE (Kingma & Welling, 2014; Rezende et al., 2014), $\beta$-VAE (Higgins et al., 2016), and $\beta$-TCVAE (Chen et al., 2018), with three aforementioned datasets. For each trained model, we manipulate one latent variable between [-3, 3] at a time while keeping others constant, and the trained decoder of the model can produce a series of images that reflect variations along that specific latent dimension. This process is repeated for each latent variable, resulting in $m\times k$ images, where $m$ is the number of latent variables and $k$ is the number of assigned values for each latent variable. We set 6 latent variables and assign the values as torch.range(-3, 3, 0.6), so $m=6$ and $k=11$ here.
Explanation generation We use GPT-4-vision as our explanation generator and compare it with a wide range of other large multimodal models: Google Bard, LLaVA-v1.5-13b (Liu et al., 2023), and InstructBLIP (Dai et al., 2024). For GPT-4-vision, LLaVA-v1.5-13b, and InstructBLIP, we all set $temperature=1,top\_p=1$. We then pass the generated images of each latent variable along with a prompt to the LLMs to produce an explanation for each latent variable. For human annotations, two annotators provide two explanations with various sentence expressions for each image sequence, so we have four human explanations for each image sequence as references. Furthermore, We evaluate the explanations across these LMMs with human annotations using BLEU(Papineni et al., 2002), ROUGE-L(Lin, 2004), METEOR(Banerjee & Lavie, 2005), and BERTScore F1(Zhang et al., 2019).
@@
lexical similarity 0.6898 0.9600 0.9412 0.9796
cosine similarity 0.9694 0.9684 1.0 0.9388
@@ 5 Case study
Figure 2 shows the explanations generated by our framework for the 3dshapes dataset. If the certainty score of the generated explanation is equal or greater than the certainty threshold ($\varepsilon=0.7434$), it will display the explanation selected. Otherwise, it will return “No clear explanation” like ${z}_{1}$ and ${z}_{5}$. Our framework can generate an explanation for each latent variable if the latent variable has an evident pattern. The explanation not only can tell what the latent variable is, but also can illustrate how the latent variable changes. Sometimes, the identified latent variable can be wrong, such as it misclassifies the wall orientation as background color in ${z}_{3}$. This is due to the deficiency in the LMM’s visual capabilities and will be elaborated in the later limitation section.
@@ 7 Conclusion
In this work, we propose a framework to comprehensively explain each latent variable in the generative models and visualize the variations of each latent variable. We first analyze the certainty scores of the explanations. The certainty score can tell us if there is a clear trend in the latent variable. We also evaluate and compare the generated explanations with various LMMs. The result shows that GPT-4-vision outperforms other large multimodal models, and we further discuss its limitations and the explanatory capabilities of various LMMs. At last, we evaluate the effects of latent variable disentanglement on the generated explanations. We believe our approach provides an efficient, explainable, and reliable way to learn the latent representations of generative models.